OpenStudy (anonymous):

Find cot theta if (-5,-12) is a point on the terminal side of theta

OpenStudy (amistre64):

i believe 5-12-13 is a righty

OpenStudy (amistre64):

thats a cot, not a cos .... i am soooo blind

OpenStudy (amistre64):

tan = y/x cot = 1/tan

OpenStudy (anonymous):

is the answer -12/13

OpenStudy (phi):